OpenStudy (anonymous):

Carissa and Louann both have savings accounts. Carissa has $250 in her account and deposits $80 per month. Louann has $1230 in her account and withdraws $60 per month. After how many months will they have equal amounts of money in their accounts? What amount of money will each have when it is the same amount? It will take months. Then, they will both have $ in their accounts.

OpenStudy (wmj259):

Equation for Carissa : $=250+80x Equation for Louann: $=1230-60x x represents the months Insert into graphing calculator or www.desmos.com Find where they intersect. Which would be at x=7 If we substitute 7 into th equations above we get $=810

OpenStudy (anonymous):

thank you

OpenStudy (wmj259):

Or instead of finding the intersection you could have set the equations equal to each other and solved for x.